Find the area under the standard normal distribution which lies between Z = -0.45 and Z = 0.75?
1
Expert's answer
2020-05-18T19:03:08-0400

"P(-0.45<Z<0.75)\\
=P(0<Z<0.45)+P(0<Z<0.75)\\
=0.1736+0.2734=0.447"